OpenStudy (anonymous):

Evaluate the expression without using a calculator. sin pi/6 + cos pi/6 I'm so lost on how to do this :/

OpenStudy (anonymous):

sin pi/6 +cos pi/6 this is same as sin 30 + cos 30

OpenStudy (dumbcow):

unit circle https://en.wikipedia.org/wiki/Unit_circle#/media/File:Unit_circle_angles_color.svg (cos,sin)

OpenStudy (anonymous):

because pi=180

OpenStudy (anonymous):

Oh okay I understand now. Thanks

OpenStudy (anonymous):

sin30= 1/2 and cos 30= \[\sqrt{3/2}\]

OpenStudy (anonymous):

now solve \[1/2 +\sqrt{3/2}\]

OpenStudy (anonymous):

It says in the back of the book that its \[1+\sqrt{3}/2\] but idk understand how they got the radical

OpenStudy (dumbcow):

from the unit circle

OpenStudy (anonymous):

so, 1+\[\sqrt{3}/2\] is the answer

OpenStudy (anonymous):

Oh I kind of see

OpenStudy (anonymous):

nevermind I do see now

OpenStudy (anonymous):

its just simple.since they have same denominator, you just add the numerators

OpenStudy (anonymous):

solve, 1/2+1/2= ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

very correct. now you understand

OpenStudy (anonymous):

thats all